# Phase diagram of generalized XY model using tensor renormalization group

### v1

@@ Abstract
We use the higher-order tensor renormalization group method to study the two-dimensional generalized XY model that admits integer and half-integer vortices. This model is the deformation of the classical XY model and has a rich phase structure consisting of nematic, ferromagnetic, and disordered phases and three transition lines belonging to the Berezinskii-Kosterlitz-Thouless and Ising class. We explore the model for a wide range of temperatures, $T$, and the deformation parameter, $\Delta$, and compute specific heat along with integer and half-integer magnetic susceptibility, finding both BKT-like and Ising-like transitions and the region where they meet.
@@ I Introduction
Spin models in two Euclidean dimensions with discrete or continuous global symmetries can have a wide range of interesting properties. Due to the famous no-go theorem known as the Hohenberg-Mermin-Wagner-Coleman (HMWC) theorem, in two-dimensional models, a continuous symmetry cannot break spontaneously, and thus, a continuous phase transition from a disordered to an ordered phase is not allowed. The simplest model in two dimensions with continuous symmetry is the classical XY model, also known as the O(2) model. In this model, the spins take values on the circle $S^{1}$. This model has been the subject of several investigations over the past fifty years [1, 2, 3, 4, 5] due to its simplicity and remarkable properties. Surprisingly, it admits an infinite-order phase transition, which is topological, known as the Berezinskii-Kosterlitz-Thouless (BKT) phase transition. This transition is peculiar since it does not follow the usual classification of the phase transitions due to Ehrenfest, which describes the order of the phase transition in terms of the lowest discontinuous derivative of the partition function. Across the phase transition in the XY model, all the derivatives of the free energy remain continuous. The transition is associated with the dissociation of the integer vortex pairs at the critical temperature. Due to the wide-ranging applications of this model, in explaining various phenomena related to superfluid Helium, thin films, superconductivity, liquid crystals, and melting of two-dimensional crystals, several extensions of this model have been considered.
One such extension we study here was first proposed by Korshunov, Lee, and Grinstein (KLG) in Refs. [6, 7]. This model has several interesting features, some of which include the possibility of fractional vortices and signs of passing directly from the disordered (high temperature) phase to the single particle (quasi) condensate phase via an Ising transition, a situation reminiscent of the ‘deconfined criticality’ scenario as studied in Ref. [8]. Due to the competition between the different terms in the Hamiltonian, this feature also leads to a richer phase structure. This model also provides a good example to understand the interplay between the discrete $\mathbb{Z}_{2}$ symmetry and $U(1)$ symmetry. The modification of the XY model, as proposed by KLG, is given by the Hamiltonian:
@@
To carry out a precise study of the phase transitions in this model, we apply the tensor network methods based on the higher-order tensor renormalization group (HOTRG) as introduced in [16]. This model has been studied using the variational uniform matrix product state (VUMPS) algorithm [17] in Ref. [18]. Here, we use an alternative approach using the numerical real-space tensor renormalization group methods. Since the proposal by Levin and Nave [19], the scheme of performing real-space coarse-graining using tensors has seen a lot of progress with extension to higher dimensions and innovative procedures to carry out the tensor renormalization  [20, 21, 22]. This method has been successfully applied to the two-dimensional XY model and related models [23, 24, 15, 25, 26, 27] and also to three-dimensional XY model with finite chemical potential [28]. Using the dual variable approach (from the character expansion), several gauge theories have also been studied [29, 30, 31] with tensors and have found agreement with results obtained using other numerical methods. The most expensive part of these computations is the contraction of the network and computation of singular values. To improve the performance of the numerical computations, we use the GPU-accelerated code described in Ref. [32].
@@ II Tensor formulation
The starting point of the real-space tensor renormalization approach is to decompose the Boltzmann weight in terms of tensor with $2d$ indices where $d$ is the number of dimensions (Euclidean). This can either be done exactly or approximately based on the symmetries of the action. One of the standard methods to do this decomposition is to use the character expansion [33], but different truncation schemes also exist [34]. Using this initial tensor one can create a network of these tensors such that when the network is contracted, it provides a good approximation to the partition function $Z$.
@@ III Numerical Results
The computation of magnetization from the first derivative of free energy (or partition function) is often prone to numerical errors. A better method is to compute the impure tensor corresponding to the magnetization and insert it into the network for the partition function [15]. To compute the dominant magnetization, depending on $\Delta$, we use either $h$ of $h_{1}$ corresponding to the different symmetry-breaking terms in the partition function. We use external magnetic field $h$ for $\Delta>0.36$ to compute $M$ and magnetic field $h_{1}$ for $\Delta\leq 0.36$ related to $\cos{(2\theta)}$, term which we refer to as $M_{1}$. We refer the reader to Appendix A for further details about the tensor construction.
Since this model has Ising and BKT-like transitions, the specific heat cannot always conclusively determine the phase transition. For such cases, we look at the first derivative of the magnetization computed using a simple finite difference method and compute the susceptibility. The peak in susceptibility signals a transition, which we then extrapolate to the zero-field limit. The extraction of this zero field critical temperature is obtained by doing functional fits of the form discussed in Ref. [15]. We show the magnetic susceptibility plot and zero-field limit extraction of the critical temperature for $\Delta=0.32$ in Fig. 2 and Fig. 3, respectively, where $\overline{h}_{1}$ is the central value used for the finite-difference of the susceptibility measurement. We refer the reader to Appendix C for plots corresponding to a wide range of $\Delta$ values. For the computation of magnetization, we use a lattice volume of $2^{35}\times 2^{35}$ while for $h=h_{1}=0$, we use a volume of $2^{30}\times 2^{30}$. For all computations, we use $D=91$ and a range of $m$ as defined in (4) to be $[-50,50]$.
@@
different $\overline{h}_{1}$ and $\Delta=0.32$.
As discussed before, the gXY model reduces to the XY model when $\Delta=1$, and much work has been done using Monte Carlo and tensor methods for the computation of the critical temperature and critical exponents. The critical temperature on the square lattice was computed to be $T_{c}=0.89290(5)$ using tensor network methods in [15], while the critical exponent $\delta=15$ was computed within errors in Ref. [32]. In this work, we start with $\Delta=0.8$, which is expected to have a single BKT-like transition, and move to smaller $\Delta$ values passing through the region where the transition lines meet the limit of $\Delta\to 0$. For the BKT transition, it is well-known that the peak of specific heat is not the correct way to determine the transition temperature (it is an overestimate, as clear from the data in Table 1). From our computation, we find that for $\Delta=0.8$, the peak of $C_{\rm{v}}$ is observed at $T=0.95(1)$. This was reported to be around $T\approx 0.91$ in Ref. [18]. To accurately determine the transition, we compute magnetic susceptibility for a small external field $h$ and then took the zero-field limit as explained above. For $\Delta=0.8$, we obtain $T\sim 0.890(4)$.
As we decrease $\Delta$ approaching the multicritical point, the difference between the transition temperatures deduced from the peak of specific heat and magnetic susceptibility, respectively, decreases for the BKT transition. As we cross the region where transition lines meet and move to smaller values of $\Delta$, we again see that the difference increases for the half-BKT line. This is evident from the data given in Table 1. Such behavior has also been noted previously in Ref. [10] where for the half-BKT line, the critical temperature values almost agree with each other around $\Delta=0.35$. We find that the specific heat peak fails to capture the correct transition temperature for $\Delta\geq 0.50$.
From the results of our tensor network computations, we find the phase diagram for this model as shown in Fig. 4. The critical temperatures for the half-BKT and BKT lines are extracted from the zero-field limit method, as stated before, and the critical temperatures for the Ising line are inferred using the peak in the specific heat. We collect the numerical results in Table 1 for critical temperature values deduced from the peak of the specific heat as well as magnetic susceptibility corresponding to the standard and nematic magnetization.
@@ IV Conclusions
We studied the phase diagram of the generalized XY model in two dimensions using the GPU-improved real-space tensor network methods. We located the phase transitions belonging to the Ising and the BKT universality classes and determined the region in the phase diagram where they appear to connect, i.e., $\Delta=0.36(2),T_{c}=0.716(3)$. Our result is consistent with an interpretation where the half-BKT and Ising transition lines first meet (as $\Delta$ is increased from zero), and then the Ising line continues for a small range of $\Delta$ to merge with the BKT line, as observed in Ref. [12]. This work refines the previous computation of this model using matrix product states (MPS) methods, and we find results that are qualitatively similar to the existing results in the literature from the standard Monte Carlo method. It remains an open problem to apply and extend the methods used in this work to models involving fractional vortices, corresponding to terms like $\cos(p\theta_{i}-p\theta_{j})$ with $p\geq 3$ in the Hamiltonian. It is believed that these models have a more complicated phase diagram, as studied in Ref. [14] using Monte Carlo methods. It would be interesting to study and revisit this model with the tensor network methods in the future.
@@ Acknowledgments
This material is based on the work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177. The work of AJ was supported in part by the Start-up Research Grant from the University of the Witwatersrand. The numerical computations were done on PARAM SMRITI, Mohali, India, supported by the Ministry of Electronics and Information Technology and Department of Science and Technology (DST), Government of India, and Syracuse University HTC Campus Grid supported by NSF award ACI-1341006. We thank Minati Biswal for collaboration during the initial stages of this work.
@@ Appendix A Derivation of tensors for the generalized O(2) model
The magnetization is calculated using (9) and is given by:
$\begin{split}M=&\;{\rm tTr}\left[\sqrt{\prod_{k=1}^{4}a_{n_{k}}(\beta,\Delta)}\;\times\;\sum_{l=-\infty}^{+\infty}I_{l}(\beta h_{1})\;\times\Bigg{(}\frac{I_{n_{1}+n_{2}-n_{3}-n_{4}+2l-1}(\beta h)+I_{n_{1}+n_{2}-n_{3}-n_{4}+2l+1}(\beta h)}{2}\Bigg{)}\right].\\ \end{split}$ (20)
We can also compute the nematic magnetization with respect to the external field $h_{1}$ and it is given as:
$\begin{split}M_{1}=&\;{\rm tTr}\left[\sqrt{\prod_{k=1}^{4}a_{n_{k}}(\beta,\Delta)}\;\times\sum_{l=-\infty}^{+\infty}I_{n_{1}+n_{2}-n_{3}-n_{4}+2l}(\beta h)\;\times\Bigg{(}\frac{I_{l+1}(\beta h_{1})+I_{l-1}(\beta h_{1})}{2}\Bigg{)}\right].\\ \end{split}$ (21)
@@ Appendix B Numerical data table
      $\Delta$       $T_{\rm{BKT};{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h},{\color[rgb]{0.78125,0,0.22265625}\definecolor[named]{pgfstrokecolor}{rgb}{0.78125,0,0.22265625}h_{1}}\to 0}$       $T_{\rm{BKT},\rm{C_{v}}}$       $T_{\rm{Ising},\rm{C_{v}}}$
@@
Table 1: The summary of the numerical results obtained in this paper. $T_{\rm{Cv}}$ is the critical temperature determined from the peak of specific heat with no external magnetic field, whereas $T_{h,h_{1}\to 0}$ is the critical temperature determined from the peak of magnetic susceptibility in the limit of vanishing external magnetic field, $h$ and $h_{1}$ respectively. The color in the third column represents the type of symmetry breaking field ($h$ for $q=1$ and $h_{1}$ for $q=2$ in  (10)) used to compute the critical temperature. For $T>0.38$, there is a single transition of the BKT universality class. Until $\Delta=0.32$, we can resolve the half-BKT and Ising line, but for $\Delta=0.34,0.36$ it is likely, based on our numerical results, that the two transition lines i.e., Ising and half-BKT have merged. For $\Delta\geq 0.40$, there is no ambiguity and transition corresponds to the BKT class. If all the transition lines meet, they do so at $\Delta=0.36(2)$. Our results are slightly more consistent with the picture that first half-BKT and Ising lines meet around $\Delta\sim 0.32$, and then the Ising line continues to merge with the BKT line around $\Delta=0.36(2)$.

### v2

@@ Abstract
We use the higher-order tensor renormalization group method to study the two-dimensional generalized XY model that admits integer and half-integer vortices. This model is the deformation of the classical XY model and has a rich phase structure consisting of nematic, ferromagnetic, and disordered phases and three transition lines belonging to the Berezinskii-Kosterlitz-Thouless (BKT) and Ising class. We explore the model for a wide range of temperatures, $T$, and the deformation parameter, $\Delta$, and compute specific heat along with integer and half-integer magnetic susceptibility, finding both BKT-like and Ising-like transitions and the region where they meet.
@@ I Introduction
Spin models in two Euclidean dimensions with discrete or continuous global symmetries can have a wide range of interesting properties. Due to the famous no-go theorem known as the Hohenberg-Mermin-Wagner-Coleman (HMWC) theorem, in two-dimensional models, a continuous symmetry cannot break spontaneously, and thus, a phase transition from a disordered to an ordered phase is not allowed. The simplest model in two dimensions with continuous symmetry is the classical XY model, also known as the O(2) model. In this model, the spins take values on the circle $S^{1}$. This model has been the subject of several investigations over the past fifty years [1, 2, 3, 4, 5] due to its simplicity and remarkable properties. Surprisingly, it admits an infinite-order phase transition, which is topological, known as the Berezinskii-Kosterlitz-Thouless (BKT) phase transition. This transition is peculiar since it does not follow the usual classification of the phase transitions due to Ehrenfest, which describes the order of the phase transition in terms of the lowest discontinuous derivative of the partition function. This transition does not violate the HMWC theorem since the transition is not due to breaking of any symmetry but due to the unbinding of vortices and antivortices (topological defects) at some finite temperature. Across the phase transition in the XY model, all the derivatives of the free energy remain continuous. The transition is associated with the dissociation of the integer vortex pairs at the critical temperature. Due to the wide-ranging applications of this model, in explaining various phenomena related to superfluid Helium, thin films, superconductivity, liquid crystals, and melting of two-dimensional crystals, several extensions of this model have been considered.
One such extension we study here was first proposed by Korshunov, Lee, and Grinstein (KLG) in Refs. [6, 7]. This model has several interesting features, some of which include the possibility of fractional vortices and signs of passing directly from the disordered (high temperature) phase to the single particle (quasi) condensate phase via an Ising transition, a situation reminiscent of the ‘deconfined criticality’ scenario as studied in Ref. [8]. Due to the competition between the different terms in the Hamiltonian, this also leads to a richer phase structure. This model also provides a good example to understand the interplay between the discrete $\mathbb{Z}_{2}$ symmetry and $U(1)$ symmetry. The modification of the XY model, as proposed by KLG, is given by the Hamiltonian:
@@
To carry out a precise study of the phase transitions in this model, we apply the tensor network methods based on the higher-order tensor renormalization group (HOTRG) as introduced in [16]. Some preliminary investigation of this model was carried out in [17]. This model has been studied using the variational uniform matrix product state (VUMPS) algorithm [18] in Ref. [19]. Here, we use an alternative approach using the numerical real-space tensor renormalization group methods. Since the proposal by Levin and Nave [20], the scheme of performing real-space coarse-graining using tensors has seen a lot of progress with extension to higher dimensions and innovative procedures to carry out the tensor renormalization  [21, 22, 23]. We refer the interested reader to Ref. [24] for summary of tensors networks based on both real-space renormalization group (RG) methods and approximation of the ground state of slightly entangled many-body systems. Tensor renormalization group (TRG) approach has been successfully applied to the two-dimensional XY model and related models [25, 26, 15, 27, 28, 29] and also to three-dimensional XY model with finite chemical potential [30]. Using the dual variable approach (from the character expansion), several gauge theories have also been studied [31, 32, 33] with tensors and have found agreement with results obtained using other numerical methods. In addition, recently some progress was also made to extract the critical behavior by studying the renormalization group collapse of magnetization for three-dimensional $O(4)$ model [34]. It seems likely that once can compute critical exponents accurately using TRG methods in coming years providing an alternative to the existing methods such as Monte Carlo and conformal bootstrap. The major part of these computations is the contraction of the network during coarse-graining and the singular value decomposition (SVD) of the growing size of tensors to restrict to fixed size after each step. This amounts to truncating the singular values to some threshold $D$ which is known as bond dimension. To improve the performance of the numerical computations in this work and get to $D=91$ in a reasonable amount of computational time, we use the grpahical processing unit (GPU)-accelerated code described in Ref. [35].
@@ II Tensor formulation
The starting point of the real-space tensor renormalization approach (on square lattice) is to decompose the Boltzmann weight in terms of tensor with $2d$ indices where $d$ is the number of dimensions (Euclidean). This can either be done exactly or approximately based on the symmetries of the action. One of the standard methods to do this decomposition is to use the character expansion [36], but different truncation schemes also exist [37]. Using this initial tensor one can create a network of these tensors such that when the network is contracted, it provides a good approximation to the partition function $Z$.
@@ III Numerical Results
The computation of magnetization from the first derivative of free energy (or partition function) is often prone to numerical errors. A better method is to compute the impure tensor corresponding to the magnetization and insert it into the network for the partition function [15]. To compute the dominant magnetization, depending on $\Delta$, we use either $h$ or $h_{1}$ corresponding to the different symmetry-breaking terms in the partition function. We use external magnetic field $h$ for $\Delta>0.36$ to compute $M$ and magnetic field $h_{1}$ for $\Delta\leq 0.36$ related to $\cos{(2\theta)}$, term which we refer to as $M_{1}$. We refer the reader to Appendix A for further details about the tensor construction.
Since this model has Ising and BKT-like transitions, the specific heat cannot always conclusively determine the phase transition. For such cases, we look at the first derivative of the magnetization computed using a simple finite difference method and compute the susceptibility. The peak in susceptibility signals a transition, which we then extrapolate to the zero-field limit. The extraction of this zero field critical temperature is obtained by doing functional fits of the form discussed in Ref. [15]. We show the magnetic susceptibility plot and zero-field limit extraction of the critical temperature for $\Delta=0.32$ in Fig. 2 and Fig. 3, respectively, where $\overline{h}_{1}$ is the central value used for the finite-difference of the susceptibility measurement. We refer the reader to Appendix C for plots corresponding to a wide range of $\Delta$ values and to Appendix D for plots corresponding to systematic error analysis when $m$ or $D$ is varied. For the computation of magnetization, we use a lattice volume of $2^{35}\times 2^{35}$ while for $h=h_{1}=0$, we use a volume of $2^{30}\times 2^{30}$. For all computations, we use $D=91$ and a range of $m$ as defined in (4) to be $[-50,50]$.
@@
different $\overline{h}_{1}$ and $\Delta=0.32$. $a=0.685(3),b=0.48(1),c=0.29(5)$.
As discussed before, the gXY model reduces to the XY model when $\Delta=1$, and much work has been done using Monte Carlo and tensor methods for the computation of the critical temperature and critical exponents. The critical temperature on a square lattice was computed to be $T_{c}=0.89290(5)$ using tensor network methods in [15], while the critical exponent $\delta\approx 15$ was computed within errors in Ref. [35]. In this work, we start with $\Delta=0.8$, which is expected to have a single BKT-like transition, and move to smaller $\Delta$ values passing through the region where the transition lines meet and going all the way down to the limit of $\Delta\to 0$. For the BKT transition, it is well-known that the peak of specific heat is not the correct way to determine the transition temperature (it is an overestimate, as clear from the data in Table 1). From our computation, we find that for $\Delta=0.8$, the peak of $C_{\rm{v}}$ is observed at $T=0.95(1)$. This was reported to be around $T\approx 0.91$ in Ref. [19]. To accurately determine the transition, we compute magnetic susceptibility for a small external field $h$ and then took the zero-field limit as explained above. For $\Delta=0.8$, we obtain $T\sim 0.890(4)$.
As we decrease $\Delta$ approaching the multicritical point (or the region where the phases meet), the difference between the transition temperatures deduced from the peak of specific heat and magnetic susceptibility, respectively, decreases for the BKT transition. As we cross the region where transition lines meet and move to smaller values of $\Delta$, we again see that the difference increases for the half-BKT line. This is evident from the data given in Table 1 in Appendix B. Such behavior has also been noted previously in Ref. [10] where for the half-BKT line, the critical temperature values almost agree with each other around $\Delta=0.35$. We find that the specific heat peak fails to capture the correct transition temperature for $\Delta\geq 0.50$.
From the results of our tensor network computations, we find the phase diagram for this model as shown in Fig. 4. The critical temperatures for the half-BKT and BKT lines are extracted from the zero-field limit method, as stated before, and the critical temperatures for the Ising line are inferred using the peak in the specific heat. We collect the numerical results in Table 1 for critical temperature values deduced from the peak of the specific heat as well as magnetic susceptibility corresponding to the standard and nematic magnetization. In Appendix C, Fig. 7-16 show the magnetic susceptibility v/s temperature plots as well as critical temperature scaling with finite external magnetic field for a variety of $\Delta$ values. In Appendix D, Table 2 we show the variation of free energy and magnetization $M_{1}$ as one varies $m$ parameter at fixed $\Delta=0.34$ near the critical temperature and Fig. 17-19 show the free energy scaling with increasing bond-dimension near the transition point different deformation parameter values.
@@ IV Conclusions
We studied the phase diagram of the generalized XY model in two dimensions using the GPU-improved real-space tensor network methods. We located the phase transitions belonging to the Ising and the BKT universality classes and determined the region in the phase diagram where they appear to connect, i.e., $\Delta=0.36(2),T_{c}=0.716(3)$. Our result is consistent with an interpretation where the half-BKT and Ising transition lines first meet (as $\Delta$ is increased from zero), and then the Ising line continues for a small range of $\Delta$ to merge with the BKT line, as observed in Ref. [12] for $q=2$ and for $q=3$ Potts model in Ref. [38]. This work refines the previous computation of this model using matrix product states (MPS) methods, and we find results that are qualitatively similar to the existing results in the literature from the standard Monte Carlo method. It remains an open problem to apply and extend the methods used in this work to models involving fractional vortices, corresponding to terms like $\cos(p\theta_{i}-p\theta_{j})$ with $p\geq 3$ in the Hamiltonian. It is believed that these models have a more complicated phase diagram, as studied in Ref. [14] using Monte Carlo methods. It would be interesting to study and revisit this model with the tensor network methods in the future.
@@ Acknowledgments
A.S. is supported by U.S. Department of Energy grant DE-SC0019139. R.G.J. is supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Co-design Center for Quantum Advantage (C2QA) under contract number DE-SC0012704 and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract number DE-AC05- 06OR23177. The work of A.J. was supported in part by the Start-up Research Grant from the University of the Witwatersrand. The numerical computations were done on PARAM SMRITI, Mohali, India, supported by the Ministry of Electronics and Information Technology and Department of Science and Technology (DST), Government of India, and Syracuse University HTC Campus Grid supported by NSF award ACI-1341006. We thank Minati Biswal for collaboration during the initial stages of this work.
@@ Appendix A Derivation of tensors for the generalized O(2) model
The magnetization is calculated using (9) and is $M=P/Z$, where $Z$ is the partition function given in (19) and $P$ is the modified contracted tensor network with impure tensor inserted. The impure tensor corresponding to $M$ is given by:
$\begin{split}\mathcal{I}_{n_{1}n_{2}n_{3}n_{4}}^{M}=&\;\left[\sqrt{\prod_{k=1}^{4}a_{n_{k}}(\beta,\Delta)}\;\times\;\sum_{l=-\infty}^{+\infty}I_{l}(\beta h_{1})\;\times\Bigg{(}\frac{I_{n_{1}+n_{2}-n_{3}-n_{4}+2l-1}(\beta h)+I_{n_{1}+n_{2}-n_{3}-n_{4}+2l+1}(\beta h)}{2}\Bigg{)}\right].\\ \end{split}$ (20)
We can also compute the nematic magnetization $M_{1}=P_{1}/Z$ with respect to the external field $h_{1}$ and its impure tensor is given as:
$\begin{split}\mathcal{I}_{n_{1}n_{2}n_{3}n_{4}}^{M_{1}}=&\;\left[\sqrt{\prod_{k=1}^{4}a_{n_{k}}(\beta,\Delta)}\;\times\sum_{l=-\infty}^{+\infty}I_{n_{1}+n_{2}-n_{3}-n_{4}+2l}(\beta h)\;\times\Bigg{(}\frac{I_{l+1}(\beta h_{1})+I_{l-1}(\beta h_{1})}{2}\Bigg{)}\right].\\ \end{split}$ (21)
@@ Appendix B Numerical data table
      $\Delta$       $T_{\rm{BKT};h,h_{1}\to 0}$       $T_{\rm{BKT},\rm{C_{v}}}$       $T_{\rm{Ising},\rm{C_{v}}}$
@@
Table 1: The summary of the numerical results obtained in this paper. $T_{\rm{Cv}}$ is the critical temperature determined from the peak of specific heat with no external magnetic field, whereas $T_{h,h_{1}\to 0}$ is the critical temperature determined from the peak of magnetic susceptibility in the limit of vanishing external magnetic field, $h$ and $h_{1}$ respectively. We use the symmetry breaking field $h$ for $q=1$ ($\Delta\in[0.38,0.80]$) and $h_{1}$ for $q=2$ ($\Delta\in[0.04,0.36]$) in Eq. (10) to compute the critical temperature.
For $T>0.38$, there is a single transition of the BKT universality class. Until $\Delta=0.32$, we can resolve the half-BKT and Ising line, but for $\Delta=0.34,0.36$ it is likely, based on our numerical results, that the two transition lines i.e., Ising and half-BKT have merged. For $\Delta\geq 0.40$, there is no ambiguity and transition corresponds to the BKT class. If all the transition lines meet, they do so at $\Delta=0.36(2)$. Our results are slightly more consistent with the picture that first half-BKT and Ising lines meet around $\Delta\sim 0.34$, and then the Ising line continues to merge with the BKT line around $\Delta=0.36(2)$.